math

use a half-angle formula to fill in blanks in the identity below: 

(cos(7x))^2= _____ + ____ cos(___)x
Answer #1

Solution:

uisng following formula to find the value of \((\cos (7 x))^{2}\)

Formula:

\(\cos ^{2}(a)=\frac{1}{2}[(1+\cos (2 a))]\)

\((\cos (7 x))^{2}=\frac{1}{2}[1+\cos (2(7 x))]\)

\(=\frac{1}{2}[1+\cos 14 x]\)

\(=\frac{1}{2}+\frac{1}{2} \cos (14 x)\)

answered by: RAMESH RAJAK